OpenStudy (anonymous):

Use an augmented matrix to solve the system of equations. Be sure to identify the row operation being used. 4x+3y=-9, x+4y=-5

OpenStudy (anonymous):

@whpalmer4 I need ya!!!

OpenStudy (e.mccormick):

Know what an augmented matrix is?

OpenStudy (anonymous):

Nope it's on my study island and we havent learned it yet :/

OpenStudy (anonymous):

well palmer I think mccormick has got me covered, but ur more then welcome to stay!

OpenStudy (e.mccormick):

if your equations were: \(ax + by = c\) and \(dx + ey = f\) then the augmented form is: \[\left[ \begin{array}{cc|c} a & b & c\\ d & e & f \end{array}\right ]\]

OpenStudy (anonymous):

ok well all I know is that u use this for something 10 01

OpenStudy (e.mccormick):

Yes. That is called the reduced, row echelon form. It gives you x and y as: \(\left[ \begin{array}{cc|c} 1 & 0 & x\\ 0 & 1 & y \end{array}\right ]\)

OpenStudy (e.mccormick):

You use elimination to get there. Add a multiple of one row to the other row.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Palmer before he left told me this If I didn't make any mistakes doing the arithmetic in my head, -21/13,-11/13 ought to be the answer (x,y)

OpenStudy (e.mccormick):

Well, put it into aug marix form and you can see.

OpenStudy (anonymous):

i know how but i dont know how to put it in to the format u have

OpenStudy (e.mccormick):

Don't worry about it looking right. Just use spaces

OpenStudy (anonymous):

k

OpenStudy (anonymous):

-21 / 13,-11 / 13 / / i dont know i used substitution in this case because one equation easily solved for a variable

OpenStudy (anonymous):

I dont think right when tired well no one does

OpenStudy (e.mccormick):

This part: 4x+3y=-9 x+4y=-5 Just use the numbers and - where needed. Remember that x means 1x.

OpenStudy (anonymous):

Thank u for ur help and effort but i have to go to bed etc. i might see u tomorrow?

OpenStudy (anonymous):

i have u fanned

OpenStudy (e.mccormick):

I should be around. Night.

OpenStudy (anonymous):

Good night to u too!